In chloroplasts at 25C, the voltage difference across the thylakoid membrane is approximately 0.03V and the pH difference across this membrane is approximately 2 pH units. What percentage of the total proton motive force in the chloroplast is due to the electrical potential difference across the thylakoid membrane? Show work. The equation we are supposed to use is pmf=Vm (membrane potential)+2.303RT *pH gradient/Faraday's constant
Answer #1

Answer

Given
Delta Vm=0.03V
pH difference (delta pH)= 2

R=8.314 X10-3 KJ
T= 250C =273+25 =298K
F= 96.485 kC/mole

Formula
pmf= Delta Vm +2.303RT *Delta pH/Faraday's constant


=(0.03)+2.303 (2) (8.314 X 10-3) kj/kg) (273+25)/96.485
=0.03 + 11.406/96485
=0.148%

In order to express in %

0.148 X 100 = 14.8%